# Impurity bands, line-nodes, and anomalous thermal Hall effect in Weyl superconductors

### v1

@@ Acknowledgments
T. Matsuhita thank Y. Masaki, M. Sato, Y. Yanase, Y. Tanaka, and J. Ieda for fruitful discussions. T. Matsushita was supported by a Japan Society for the Promotion of Science (JSPS) Fellowship for Young Scientists and by JSPS KAKENHI Grant No. JP19J20144 and JST CREST Grant No. JPMJCR19T2. This work was also supported by JST CREST Grant No. JPMJCR19T5, Japan, and JSPS KAKENHI (Grants No. JP21H01039 and No. JP22H01221). I. V. was supported in part by grant NSF PHY-1748958 to the Kavli Institute for Theoretical Physics (KITP).
@@
In this appendix, we derive the linear response of anomalous Keldysh Green function, $\delta\check{g}^{a}$, to a temperature gradient. We start from the Eilenberger equation for the non-equilibrium Keldysh Green function, $\delta\check{g}^{\rm K}$,
@@
With the anomalous Keldysh Green function, $\delta\underline{g}^{a}$, and the anomalous Keldysh impurity self-energy, $\delta\underline{\sigma}_{\rm imp}^{a}$, we recast Eq. (A) into,
@@
The anomalous Keldysh impurity self-energy is calculated from the $T$-matrix equation,
@@
The anomalous Keldysh Green function satisfies $\underline{g}_{\rm eq}^{\rm R}\delta\underline{g}^{a}+\delta\underline{g}^{a}\underline{g}_{\rm eq}^{\rm A}=0$. Using this normalization, we can solve the transport equation, (A), and obtain the anomalous Keldysh Green function,
@@
At low temperature, we expand the Green function in the energy $\epsilon$, because we expect the energy range to be cut off by the temperature, while the Green function varies on the scale of the superconducting gap. As seen in the main text, the non-equilibrium retarded and advanced functions do not contribute to thermal transport because they arise from the temperature dependence of the equilibrium gap function. We thus focus on the anomalous Keldysh Green function. To the leading order in $\epsilon$, the anomalous Keldysh Green function, Eqs. (23a)-(23c) reduce to,
@@
×$\times\left(\underline{g}_{\rm eq,LT}^{\rm R}\delta\underline{\sigma}_{\rm imp,LT}^{a}-\delta\underline{\sigma}_{\rm imp,LT}^{a}\underline{g}_{\rm eq,LT}^{\rm A}\right).$ (41)
Here the subscript “${\rm LT}$” denotes the zero-frequency limit, $\epsilon=0$. In that limit, the impurity self-energy $\sigma_{\rm imp,eq0}$ in equilibrium is purely real and the retarded and advanced values are identical. $\underline{g}^{\rm R(A)}_{\rm eq,LT}$ is given by,
@@
with $D_{\rm LT}\equiv D^{\rm R}_{\rm LT}=D^{\rm A}_{\rm LT}=\sqrt{\gamma^{2}+|\Delta_{\rm eq}\hat{k}_{{\rm F}\perp}\eta(\hat{k}_{z})|^{2}}$. The self-energy $\sigma_{\rm imp,LT}^{\rm R}=-i\gamma\underline{\tau}_{z}$ in equilibrium is calculated from the $T$-matrix equation,
@@
Here, we assume a spatially uniform temperature gradient along the $y$-direction and $\Delta_{\rm eq}\in\mathbb{R}$. As in the manuscript, we consider the $d$-vector fixed along the $z$ axis. In this case, the spin along the $z$ axis is conserved, which allows us to focus on each spin subspace. Hence, we perform the low temperature expansion in each spin subspace and drop the spin index. In the zero-frequency limit, the anomalous Keldysh impurity self-energy is calculated from the self-consistent $T$-matrix equation,
@@
From the Eq. (B), we make an ansatz for the anomalous Keldysh impurity self-energy,
@@
With this anomalous Keldysh impurity self-energy, we transform the $T$-matrix equation (B) into,
@@
From the matrix equation (49), we obtain the coefficient for the anomalous Keldysh impurity self-energy, $\tilde{X},\;\tilde{Y}$,
@@
We now obtain the low temperature formula for thermal conductivities,
@@
In the clean system $\Gamma_{\rm imp}\ll\pi T_{c}$, the low temperature formula for the thermal conductivity reduces to Eqs. (25) and (26),
@@
We give the derivation of Eq. (30) to clarify the relation between the extrinsic ATHE and the formation of the impurity band. To associate these, we consider the equilibrium $T$-matrix equation and the zero-energy DOS. Differentiating Eqs. (44) and (45) with the scattering phase-shift, we obtain

### v2

@@ Acknowledgments
T. Matsuhita thank Y. Masaki, M. Sato, Y. Yanase, Y. Tanaka, J. Ieda, H. Kontani, S. Onari, A. Kobayashi, and R. Tazai for fruitful discussions. T. Matsushita was supported by a Japan Society for the Promotion of Science (JSPS) Fellowship for Young Scientists, JSPS KAKENHI Grant No. JP19J20144, No. 24KJ0130, and JST CREST Grant No. JPMJCR19T2. This work was also supported by JST CREST Grant No. JPMJCR19T5, Japan, and JSPS KAKENHI (Grants No. JP21H01039 and No. JP22H01221). I. V. was supported in part by grant NSF PHY-1748958 to the Kavli Institute for Theoretical Physics (KITP).
@@
In this appendix, we derive the anomalous Keldysh QGF, $\delta\check{g}^{a}$, as the linear response to temperature gradients. We start from the Eilenberger equation for the nonequilibrium Keldysh QGF, $\delta\check{g}^{\rm K}$,
@@
Using the anomalous Keldysh QGF and $\delta\underline{\sigma}_{\rm imp}^{a}$, we recast Eq. (A) into,
@@
$\delta\underline{\sigma}_{\rm imp}^{a}$ is calculated from the $T$-matrix equation,
@@
The anomalous Keldysh QGF is supplemented by $\underline{g}_{\rm eq}^{\rm R}\delta\underline{g}^{a}+\delta\underline{g}^{a}\underline{g}_{\rm eq}^{\rm A}=0$ from the normalization condition. Using this condition, we can solve the transport equation [Eq (A)] and obtain the anomalous Keldysh QGF,
@@
At low temperatures, we can expand the QGF in the frequency $\epsilon$ because the anomalous Keldysh QGF is proportional to the derivative of the equilibrium distribution function, which introduces a small cut-off in the frequency integral in Eq. (16). As described in the main text, the change in the spectrum function from equilibrium does not contribute thermal transport. Therefore, we focus on the anomalous Keldysh QGF.
To the leading order in $\epsilon$, the anomalous Keldysh QGF [Eq. (23)] reduces to,
@@
×$\times\left(\underline{g}_{\rm eq,LT}^{\rm R}\delta\underline{\sigma}_{\rm imp,LT}^{a}-\delta\underline{\sigma}_{\rm imp,LT}^{a}\underline{g}_{\rm eq,LT}^{\rm A}\right),$ (44)
where the subscript “${\rm LT}$” represents the zero-frequency limit, $\epsilon=0$. $\underline{g}^{\rm R,A}_{\rm eq,LT}$ is the equilibrium QGF in this limit, which is given by,
@@
with $D_{\rm LT}\equiv D^{\rm R}_{\rm LT}=D^{\rm A}_{\rm LT}=\sqrt{\gamma^{2}+|\Delta_{\rm eq}\hat{k}_{{\rm F}\perp}\eta(\hat{k}_{z})|^{2}}$. In the zero-frequency limit, $\sigma_{\rm imp,eq0}$ is purely real and irrelevant for the thermal transport at low-temperatures. Thus, the equilibrium self-energy expressed as $\underline{\sigma}_{\rm imp,LT}^{\rm R}=-i\gamma\underline{\tau}_{z}$, where $\sigma_{\rm imp,eq0}$ is neglected. $\gamma$ is calculated from the $T$-matrix equation,
@@
Here, we consider a spatially uniform temperature gradient along the $y$-direction. We also assumes $\Delta_{\rm eq}\in\mathbb{R}$ without loss of generality. In this paper, we consider the $d$-vector fixed along the $z$ axis. In this case, the spin along the $z$ axis is conserved, which allows us to focus on each spin subspace. Hence, we perform the low-temperature expansion in each spin subspace and drop the spin index. In the zero-frequency limit, $\delta\underline{\sigma}_{\rm imp}^{a}$ is calculated from the self-consistent $T$-matrix equation,
@@
From the Eq. (B), we make an ansatz for $\delta\underline{\sigma}_{\rm imp,LT}^{a}$,
@@
Using this ansatz, we transform the $T$-matrix equation [Eq. (B)] into,
@@
From this matrix equation, we obtain the coefficient for $\delta\underline{\sigma}_{\rm imp,LT}^{a}$,
@@
We finally obtain the low-temperature formula for thermal conductivities,
@@
In the clean system $\Gamma_{\rm imp}\ll\pi T_{c}$, the low-temperature formula for the thermal conductivity reduces to Eqs. (27) and (28),
@@
We provide the derivation of Eq. (32) to establish the relationship between the extrinsic ATHE and the change in the DOS due to impurity bands. We consider the equilibrium $T$-matrix equation and the zero-energy DOS. Differentiating Eqs. (47) and (48) with the scattering phase-shift, we obtain